# Pointer-Generator Networks for Low-Resource Machine Translation: Don’t Copy That!

### v1

@@ Abstract
While Transformer-based neural machine translation (NMT) is very effective in high-resource settings, many languages lack the necessary large parallel corpora to benefit from it. In the context of low-resource (LR) MT between two closely-related languages, a natural intuition is to seek benefits from structural “shortcuts”, such as copying subwords from the source to the target, given that such language pairs often share a considerable number of identical words, cognates, and borrowings. We test Pointer-Generator Networks for this purpose for six language pairs over a variety of resource ranges, and find weak improvements for most settings. However, analysis shows that the model does not show greater improvements for closely-related vs. more distant language pairs, or for lower resource ranges, and that the models do not exhibit the expected usage of the mechanism for shared subwords. Our discussion of the reasons for this behaviour highlights several general challenges for LR NMT, such as modern tokenization strategies, noisy real-world conditions, and linguistic complexities. We call for better scrutiny of linguistically motivated improvements to NMT given the blackbox nature of Transformer models, as well as for a focus on the above problems in the field.
@@ 1 Introduction and Motivation
While state-of-the-art Transformer models (Vaswani et al., 2017) for NMT work well for high-resource language pairs, their performance degrades in low-resource situations (Kim et al., 2020; Haddow et al., 2022); this means that most languages in the world cannot benefit from mainstream advances and models (Joshi et al., 2020). There is therefore a clear appeal to developing simple architectural mechanisms for these models that are targeted at yielding improvements in data-scarce scenarios, while interfering minimally with mainstream preprocessing, tokenization, and training pipelines.
In the context of a low-resource language (LRL), we are often interested in translation to and from a closely related HRL, which possible has linguistic genealogical, regional, and cultural ties with the LRL, [^1] This is the case, for example, for several dialects of Arabic, all closely related to relatively high-resource Modern Standard Arabic, and languages of the Turkic and Indic language continua. in order to make the abundant content in HRLs available in related LRLs. We expect that closely related languages share considerable overlap at the subword level from cognates, borrowings and shared vocabulary (see examples in Figure 1). Given the absence of large parallel corpora for our language pair, we aim to leverage this shared knowledge across source and target, intuitively, to provide “easier” routes for our MT model from source to target sentence.
@@
Pointer Generator Networks (PGNs; See et al. (2017)) are a mechanism which allow the model, for every output token produced, to either copy some token from the input (“point”) or “generate” a token as per usual from the vocabulary. PGNs have been used for a variety of problems, described in Section 2; however, as we far as we know, this is the first work to study its applicability to LR NMT.
We introduce a PGN mechanism into a Transformer-based NMT architecture, and test its performance for $6$ language pairs over $4$ low-resource ranges. Hindi-Bhojpuri (hi-bh), Spanish-Catalan (es-ca), French-Occitan (fr-oc) represent closely-related pairs, and Hindi-Marathi (hi-mr), Spanish-English (es-en), and French-German (fr-de) represent more distant pairs. We expect that PGN will help most for (1) lower-resource scenarios (2) more closely-related language pairs (3) sentence pairs with higher subword overlap. While PGN shows improvements in certain settings, our analysis shows that they do not come from intended sources; in particular, we find clear lack of evidence for the above hypotheses. We discuss various factors that contribute to this failure, highlighting fundamental challenges for LR NMT, such as noisy datasets, mainstream tokenization practices best suited for high-resource scenarios, as well as linguistic and orthographic complexities that may obfuscate underlying source-target similarities.  [^2] Code will be made public in non-anonymous version.
@@ 2 Related Work
Pointer Networks were introduced to solve problems that involved permuting the input, such as the Traveling Salesman Problem and the complex hull problem (Vinyals et al., 2015). Their use in NLP has been largely been for monolingual summarization, where the target may naturally contain identical spans from the source. Cheng and Lapata (2016) present a complex hierarchical LSTM-based model for summarization, which directly extracts sentences from the text and words from sentences. Gulcehre et al. (2016) use pointer networks in RNN-based sequence-to-sequence models for summarization and machine translation, training their model explicitly to use the pointing mechnanism for uncommon words. Gu et al. (2016) and See et al. (2017) also incorporate variants of pointer-generator networks into RNN-based sequence-to-sequence learning for summarization. Prabhu and Kann (2020) applied PGNs to the task of grapheme-to-phoneme conversion via an explicit source-target mapping. Zhang et al. (2021) proposed a pointer-disambiguator-copier (PDC) system for dictionary-enhanced high-resource NMT, using source word translations as potential candidates for the copying mechanism, with a disambiguator component to select appropriate senses.
@@ 3 Model
Here, $\textbf{c}^{t}$ is the context vector, calculated as $\textbf{c}^{t}=(\textbf{a}^{t})^{T}\textbf{e}^{t}$, where $\textbf{a}^{t}$ represents cross-attention vector, and $\textbf{e}^{t}$ contains the encoder hidden states. $\textbf{d}^{t}$ and $\textbf{x}^{t}$ contain the decoder’s final hidden states and input respectively, $\oplus$ denotes concatenation, and W and B are learned weights and a bias vector respectively. $\textbf{P}_{\textbf{c}}^{t}$ and $\textbf{P}_{\textbf{g}}^{t}$ represent the copy and generate distributions (softmaxed logits) respectively at step $t$. We use cross-attentions weights over source tokens for the copy logits, and standard decoder outputs for generate logits.
@@ 4 Experiments
We used the WikiMatrix (wm) corpus (Schwenk et al., 2019) for es-en, es-ca, fr-de and fr-oc. For es-ca, we also report results on synthetic Europarl (ep) parallel data (Koehn, 2005). [^3] The Europarl dataset was automatically translated into Catalan; taken from https://github.com/Softcatala/Europarl-catalan. For hi-mr, we used the CVIT-PIB corpus (Philip et al., 2021), and for the low-resource pair hi-bh, we use the NLLB corpus (Schwenk et al., 2021; Team et al., 2022; Heffernan et al., 2022). See Table 1 for dataset heuristics. Note the higher per token overlap as expected for our closely-related group as compare to the others. The hi-bh sentences are extremely short, and share fewer tokens than expected: in this case, this reflects badly parallel data. [^4] See App A for more details on datasets.
@@ 5 Results and Discussion
We record values of $p_{copy}$ to track the model’s usage of the copy mechanism. While $p_{copy}$ values are relatively high [^7] Note that it is difficult to comment on absolute values of $p_{copy}$. The copying distribution is normalized over the sentence length whereas generate distributions are normalized over the vocabulary; even low values of $p_{copy}$ will considerably affect the mixed distribution. for copied subwords, numerals, and proper nouns, we often see that they they are also high for seemingly random subwords. [^8] See App C for visualizations of this behaviour. Examples with counter-intuitively high values of $p_{copy}$: quiero-vull (es-ca), behad-atishay (hi-mr). We also do not see a relationship between the $p_{copy}$ value of a target token and the entropy of the cross-attention distribution for that token.
A reasonable intuition about PGN training generalization is that in the absence of any information, the model will default to copying, since this is likely to do better on average than guesses over the entire vocabulary, and that eventually, it will learn to generate language-specific subwords, memorising the relevant strategy for given subwords in encoder states (used to calculate $p_{copy}$ as shown in Section 3). However, our visualizations of cross-attention and $p_{copy}$ usage throughout training show no evidence of this generalization strategy. It’s possible that since initial cross-attention distributions are noisy, and most subwords are not direct copies, the model is discouraged from copying early on; it’s also possible that the model finds it easier or trivial to encode copied source-target equivalents via the “generate” mechanism and does not need an explicit copier, given that it must additionally learn which subwords should be copied. In general, it appears that the model uses the copy mechanism to encode a task that is not easily interpretable. We discuss potential reasons for this below.
@@
In theory, the copier would learn best if the tokenizer behaved in a morphologically principled manner.999e.g. given khaya-khalla (ate) in Hindi and Marathi, we ideally want kha ##ya and kha ##lla. This will allow the common stem kha to be copied over, while the language specific inflection subwords can be generated. However, BPE tokenization generally results in subword splits that may not reflect shared stems in word equivalents.101010e.g. our trained tokenizer contains both propuesto (es) and proposat (ca) instead of sharing the subword prop.
A natural idea here may be an investigation of morphologically inspired tokenizers (Pan et al., 2020). However, we generally see inconclusive, at best marginal, benefits of such tokenizers over BPE in modern neural MT (Macháček et al., 2018; Domingo et al., 2019; Mielke et al., 2021), especially those relying on unsupervised mophological segmentation, e.g., with Morfessor (Creutz and Lagus, 2007) in the absence of morphological analysers; these ideas have not been incorporated into mainstream tokenization strategies.
@@
While closely-related language show high (subword) vocabulary overlap, word equivalences may be obscured by sound change and orthographic systems; if these changes are word-internal, then even an ideal tokenizer will see different stems/tokens in the source and target.111111e.g. vishwas-biswas (hi-bh, sound change), website-Webseite (en-de, orthographic system) Further, we may see that a word that has a cognate in its sister language is translated to a non-cognate, due to semantic drift, or differences in idiom or usage norms in the two languages, e.g. kitaab-pustak (hi-mr), resulting in non-identical subword equivalences. These phenomena are often unpredictable and unsystematic; even if not, they are not trivial to model into tokenization or architectural strategies for MT. [^12] See App B for experiments with minor variants of our approach dealing with pretrained encoder/decoder initialization, tokenizer size, choice of attention head, and identical source-target settings.
@@ 6 Challenges for LR NMT
Incorporating knowledge of linguistic relationships among closely related data-imbalanced language pairs offers a natural strategy for mitigating data scarcity in mainstream NMT between regional languages, and it is crucial to understand the challenges in this realm. We show that while the PGN mechanism offers an intuitive theoretical shortcut for translation between closely related languages, its performance in practice is limited, potentially by the combined effect of noisy real-world datasets containing non-literal translations, the behaviour of standard tokenisers, as well as linguistic complexities beyond simplified ideas of shared vocabulary and cognates. These are inevitable hurdles to any project that attempts to use structural linguistic knowledge to benefit NMT performance.
@@
Finally, our analysis hints that PGN-like shortcuts may not be worth offering in the first place: “easy” equivalences, a natural target of linguistic interventions, may not be the bottleneck for LR NMT. Instead, it is more likely that the true bottlenecks are handling precisely the above challenges, i.e. non-systemic differences, one-off phenomena, and real-world noise in low-resource conditions.
@@ 7 Limitations
Further, our results are limited to the $6$ language pairs that we experimented on. While we simulate identical low-resource conditions for all our language pairs, we clearly see the difference in absolute performances on hi-mr or hi-bh as compared to the high-resource language pairs: the data for the latter are simply of much better quality. This demonstrates the need to experiment and present results on truly low-resource conditions: we include hi-bh as such a language pair. Finally, this discussion is only relevant to translation between closely-related languages that share a script (although this is the predominant case), allowing for lexical similarity to be reflected by shared subwords.
@@ Appendix C Visualizations
Note that the hi-bh source-target sentence pairs are not in fact translations of each other and exemplify the noise we discuss in Section 4 and App A. The cross-attention distributions for the es-ca and fr-oc models are in general much better defined and able to attend to appropriate tokens (in these example visualizations as well as others that we looked at); this is a consequence of the better quality of the data and models in these languages.

### v3

@@ Abstract
While Transformer-based neural machine translation (NMT) is very effective in high-resource settings, many languages lack the necessary large parallel corpora to benefit from it. In the context of low-resource (LR) MT between two closely-related languages, a natural intuition is to seek benefits from structural “shortcuts”, such as copying subwords from the source to the target, given that such language pairs often share a considerable number of identical words, cognates, and borrowings. We test Pointer-Generator Networks for this purpose for six language pairs over a variety of resource ranges, and find weak improvements for most settings ($<1$ BLEU). However, analysis shows that PGNs do not show greater improvements for closely-related vs. more distant language pairs, or for lower resource ranges, and that the models do not exhibit the expected usage of the mechanism for shared subwords. Our discussion of the reasons for this behaviour highlights several general challenges for LR NMT, such as modern tokenization strategies, noisy real-world conditions, and linguistic complexities. We call for better scrutiny of linguistically motivated improvements to NMT given the blackbox nature of Transformer models, as well as for a focus on the above problems in the field.
@@ 1 Introduction and Motivation
While state-of-the-art (SOTA) Transformer models (Vaswani et al., 2017) for NMT work well for high-resource language pairs, their performance degrades in low-resource situations (Koehn and Knowles, 2017; Sennrich and Zhang, 2019; Kim et al., 2020; Haddow et al., 2022); this means that most languages in the world cannot benefit from mainstream advances and models (Joshi et al., 2020). There is therefore a clear appeal to developing simple architectural mechanisms for these models that are targeted at yielding improvements in data-scarce scenarios, while interfering minimally with mainstream preprocessing, tokenization, and training pipelines.
In the context of a low-resource language (LRL), we are often interested in translation to and from a closely related HRL, which possibly has linguistic genealogical, regional, and cultural ties with the LRL, [^1] This is the case, for example, for several languages of the Arabic continuum, all closely related to relatively high-resource Modern Standard Arabic, and languages of the Turkic and Indic language continua. in order to make the abundant content in HRLs available in related LRLs. We expect that closely related languages share considerable overlap at the subword level from cognates, borrowings and shared vocabulary (see examples in Figure 1). Given the absence of large parallel corpora for our language pair, we aim to leverage this shared knowledge across source and target, intuitively, to provide “easier” routes for our MT model from source to target sentence.
@@
Pointer Generator Networks (PGNs; See et al. (2017)) are a mechanism which allow the model, for every output token produced, to either copy some token from the input (“point”) or “generate” a token as per usual from the vocabulary. PGNs have been used for a variety of problems, described in Section 2, often targeted at repeated spans of text in the input and output; however, as we far as we know, this is the first work to study its applicability to LR NMT. In this case, we hypothesise that the pointing mechanism will show advantages for rare shared subwords, for which the best strategy may be to copy them to the output.
We introduce a PGN mechanism into a Transformer-based NMT architecture, and test its performance for 6 language pairs over 4 low-resource training ranges. We work with Hindi-Bhojpuri (hi-bh), Spanish-Catalan (es-ca), and French-Occitan (fr-oc), representing closely-related pairs, Hindi-Marathi (hi-mr), a relatively more distant pair, [^2] Hindi, Bhojpuri, and Marathi belong to the Indic branch of the Indo-European family. Hindi and Bhojpuri further belong to the Shaurasenic sub-branch and are closer lexically and grammatically to each other and other languages on or close to the Hindi Belt such as Punjabi, Rajasthani, Haryanvi, and Maithili, than Hindi is to Marathic languages and dialects; this is supported by lexical and other studies of cross-lingual similarity (Sengupta and Saha, 2015; Mundotiya et al., 2021; Bafna et al., 2022). See Glottolog (https://glottolog.org/resource/languoid/id/cont1248) for the phylogenetic tree. and Spanish-English (es-en) and French-German (fr-de), representing further distant pairs. We expect that PGN will help most for (1) lower-resource scenarios (2) more closely-related language pairs (3) sentence pairs with higher subword overlap. While PGN shows improvements in certain settings, our comparative analysis of the benefits of PGNs across the three dimensions above shows clear lack of evidence for the these hypotheses. Further, our visualizations of the PGN mechanism also indicate that observed benefits do not come from intended sources. We discuss various factors that contribute to this failure, highlighting fundamental challenges for LR NMT, such as noisy datasets, mainstream tokenization practices best suited for high-resource scenarios, as well as linguistic and orthographic complexities that may obfuscate underlying source-target similarities. [^3] https://github.com/niyatibafna/pgns-for-lrmt
@@ 2 Related Work
Pointer Networks were introduced to solve problems that involved permuting the input, such as the Traveling Salesman Problem and the complex hull problem (Vinyals et al., 2015). Their use in NLP has been largely been for monolingual summarization, where the target may naturally contain identical spans from the source. Cheng and Lapata (2016) present a complex hierarchical LSTM-based model for summarization, which directly extracts sentences from the text and words from sentences. Gulcehre et al. (2016) use pointer networks in RNN-based sequence-to-sequence models for summarization and machine translation, training their model explicitly to use the pointing mechanism for uncommon words. Gu et al. (2016) and See et al. (2017) also incorporate variants of pointer-generator networks into RNN-based sequence-to-sequence learning for summarization. Prabhu and Kann (2020) applied PGNs to the task of grapheme-to-phoneme conversion via an explicit source-target mapping. Zhang et al. (2021) proposed a pointer-disambiguator-copier (PDC) system for dictionary-enhanced high-resource NMT, using source word translations as potential candidates for the copying mechanism, with a disambiguator component to select appropriate senses.
@@ 3 Model
Here, $\textbf{c}^{t}$ is the context vector, calculated as $\textbf{c}^{t}=(\textbf{a}^{t})^{T}\textbf{e}^{t}$, where $\textbf{a}^{t}$ represents cross-attention vector, and $\textbf{e}^{t}$ contains the encoder hidden states. $\textbf{d}^{t}$ and $\textbf{s}^{t}$ contain the decoder’s final hidden states and input respectively, $\oplus$ denotes concatenation, and W and B are learned weights and a bias vector respectively. $\textbf{P}_{\textbf{c}}^{t}$ and $\textbf{P}_{\textbf{g}}^{t}$ represent the copy and generate distributions (softmaxed logits) respectively at step $t$. We use cross-attention weights over source tokens for the copy logits, and standard decoder outputs for generate logits.
@@ 4 Experiments
We used the WikiMatrix (wm) corpus (Schwenk et al., 2019) for es-en, es-ca, fr-de and fr-oc. For es-ca, we also report results on synthetic Europarl (ep) parallel data (Koehn, 2005). [^4] The Europarl dataset was automatically translated into Catalan; taken from https://github.com/Softcatala/Europarl-catalan. For hi-mr, we used the CVIT-PIB corpus (Philip et al., 2021), and for the low-resource pair hi-bh, we use the NLLB corpus (Schwenk et al., 2021; Team et al., 2022; Heffernan et al., 2022). See Table 1 for dataset heuristics. Note the higher per token overlap as expected for our closely-related group as compared to the others. The hi-bh sentences are extremely short, and share fewer tokens than expected: in this case, this reflects badly parallel data. [^5] See Appendix A for more details on datasets.
@@ 5 Results and Discussion
We record values of $p_{copy}$ to track the model’s usage of the copy mechanism. While $p_{copy}$ values are relatively high [^8] Note that it is difficult to comment on absolute values of $p_{copy}$. The copying distribution is normalized over the sentence length whereas generate distributions are normalized over the vocabulary; even low values of $p_{copy}$ will considerably affect the mixed distribution. for copied subwords, numerals, and proper nouns, we often see that they they are also high for seemingly random subwords. [^9] See Appendix C for visualizations of this behaviour. Examples with counter-intuitively high values of $p_{copy}$: quiero-vull (es-ca), behad-atishay (hi-mr). We also do not see a relationship between the $p_{copy}$ value of a target token and the entropy of the cross-attention distribution for that token.
A reasonable intuition about PGN training generalization is that in the absence of any information, the model will default to copying, since this is likely to do better on average than guesses over the entire vocabulary, and that eventually, it will learn to generate language-specific subwords, memorising the relevant strategy for given subwords in encoder states (used to calculate $p_{copy}$ as shown in Section 3). However, our visualizations of cross-attention and $p_{copy}$ usage throughout training show no evidence of this generalization strategy. It’s possible that since initial cross-attention distributions are noisy, and most subwords are not direct copies, the model is discouraged from copying early on; it’s also possible that the model finds it easier or trivial to encode copied source-target equivalents via the “generate” mechanism and does not need an explicit copier, given that it must additionally learn which subwords should be copied. We discuss potential reasons for this below. In general, it appears that the model uses the copy mechanism to encode a task that is not easily interpretable, possibly resulting in the observed small improvements over some datasets.
@@
In theory, the copier would learn best if the tokenizer behaved in a morphologically principled manner.101010e.g. given khaya-khalla (ate) in Hindi and Marathi, we ideally want kha ##ya and kha ##lla. This will allow the common stem kha to be copied over, while the language specific inflection subwords can be generated. However, BPE tokenization generally results in subword splits that may not reflect shared stems in word equivalents (Ataman and Federico, 2018).111111e.g. our trained tokenizer contains both propuesto (es) and proposat (ca) instead of sharing the subword prop.
A natural idea here may be an investigation of morphologically inspired tokenizers (Pan et al., 2020; Ortega et al., 2020; Chen and Fazio, 2021). However, we generally see inconclusive, at best marginal, benefits of such tokenizers over BPE in modern neural MT (Macháček et al., 2018; Domingo et al., 2019; Mielke et al., 2021), especially those relying on unsupervised morphological segmentation, e.g., with Morfessor (Creutz and Lagus, 2007) in the absence of morphological analysers. These ideas have not been incorporated into mainstream tokenization strategies.
@@
While closely-related language show high (subword) vocabulary overlap, word equivalences may be obscured by sound change and orthographic systems; if these changes are word-internal, then even an ideal tokenizer will see different stems/tokens in the source and target.121212e.g. vishwas-biswas (hi-bh, sound change), website-Webseite (en-de, orthographic system) Further, we may see that a word that has a cognate in its sister language is translated to a non-cognate, due to semantic drift, or differences in idiom or usage norms in the two languages, e.g. kitaab-pustak (hi-mr), resulting in non-identical subword equivalences. These phenomena are often unpredictable and unsystematic; even if not, they are not trivial to model into tokenization or architectural strategies for MT.
See Appendix B for experiments with minor variants of our approach dealing with pretrained encoder/decoder initialization, tokenizer size, choice of attention head, and identical source-target settings.
@@ 6 Challenges for LR NMT
Incorporating knowledge of linguistic relationships among closely related data-imbalanced language pairs offers a natural strategy for mitigating data scarcity in mainstream NMT between regional languages, and it is crucial to understand the challenges in this realm. We show that while the PGN mechanism offers an intuitive theoretical shortcut for translation between closely related languages, its performance in practice is limited, potentially by the combined effect of noisy real-world datasets containing non-literal translations, the behaviour of standard tokenizers, as well as linguistic complexities beyond simplified ideas of shared vocabulary and cognates. These are inevitable hurdles to any project that attempts to use structural linguistic knowledge to benefit NMT performance.
@@
Finally, our analysis hints that PGN-like shortcuts may not be worth offering in the first place: “easy” equivalences, a natural target of linguistic interventions, may not be the bottleneck for LR NMT. Instead, it is more likely that the true bottlenecks are handling precisely the above challenges, i.e. non-systematic differences, one-off phenomena, and real-world noise in low-resource conditions.
@@ 7 Limitations
Further, our results are limited to the $6$ language pairs that we experimented on. While we simulate identical low-resource conditions for all our language pairs, we clearly see the difference in absolute performances on hi-mr or hi-bh as compared to the high-resource language pairs: the data for the latter are simply of much better quality. This demonstrates the need to experiment and present further results on non-simulated truly low-resource conditions, such as the hi-bh language pair studied here. Finally, this discussion is only relevant to translation between closely-related languages that share a script (although this is the predominant case), allowing for lexical similarity to be reflected by shared subwords.
8 Conclusion
In this work, we investigate the applicability of Pointer-Generator Networks in NMT, hypothesizing that an explicit copy mechanism will provide benefits for low-resource translation between closely related languages. We show that while we do observe weak improvements, these are not higher for closer-related languages, sentence pairs with higher overlap, or lower resource ranges, contrary to intuition. Our discussion of potential reasons for the failure of this approach highlights several general challenges for low-resource NMT, such as mainstream tokenization strategies, noisy data, and non-systematic linguistic differences.
@@ Appendix C Visualizations
Note that the hi-bh source-target sentence pairs are not in fact translations of each other and exemplify the noise we discuss in Section 4 and Appendix A. The cross-attention distributions for the es-ca and fr-oc models are in general much better defined and able to attend to appropriate tokens (in these example visualizations as well as others that we looked at); this is a consequence of the better quality of the data and models in these languages.